OpenStudy (anonymous):

write the solution in interval notation :) 3/5x-2 < 3/10-x

OpenStudy (helder_edwin):

\[ \large \frac{3}{5x-2}<\frac{3}{10-x} \] ??

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 } x -2 < \frac{ 3 }{ 10 } - x\]

OpenStudy (anonymous):

@helder_edwin

OpenStudy (helder_edwin):

way easier then.

OpenStudy (helder_edwin):

put the x's on one side (LHS) and the numbers on the other side (RHS)

OpenStudy (anonymous):

not sure how to do that

OpenStudy (helder_edwin):

i mean add x and 2 to both sides

OpenStudy (helder_edwin):